OpenStudy (anonymous):

Let f(x)=2x+2 solve f^-1(x)when x=4

myininaya (myininaya):

basically to find f^(-1)(x) you could actually find that formula for f^(-1)(x) or since f^(-1)(4)=y implies f(y)=4 you could just solve f(y)=4 for y to get your answer so f(y)=2y+2 f(y)=4 replace f(y) with 2y+2 so we have 2y+2=4

myininaya (myininaya):

solve